8-z/3 > or = 11 how do you work this
Answer by Adrish(44)

Solutions:
8 - z/3>= 11
8 -8 -z/3 >= 11-8
-z/3 >= 3
Multiply both sides by 3 we get
-z >= 9
z<=-9